Phone calls arrive at the rate of 48 per hour at the reservation desk for Regional Airways.

a. Compute the probability of receiving two calls in a 10-minute interval of time.

(to 4 decimals)

b. Compute the probability of receiving exactly 13 calls in 15 minutes.

(to 4 decimals)

c. Suppose no calls are currently on hold. If the agent takes 5 minutes to complete the current call, how many callers do you expect to be waiting by that time?

What is the probability that none will be waiting?

(to 4 decimals)

d. If no calls are currently being processed, what is the probability that the agent can take 1 minutes for a personal time without being interrupted by a call?

(to 4 decimals)
Answer #1

a)expected number of calls in 10 minute=48*10/60=8

probability of receiving two calls in a 10-minute interval of time =e-8*82/2! =0.0107

b)

probability of receiving exactly 13 calls in 15 minutes =e-12*1213/13! =0.1056

c) expect to be waiting by that time=5*48/60 =4

d) probability that the agent can take 1 minutes for a personal time without being interrupted by a call =0.4493